OpenStudy (anonymous):

What is the sum of a 30-term arithmetic sequence where the first term is 72 and the last term is -102?

jimthompson5910 (jim_thompson5910):

Hint: The formula is S = n*(first term + last term)/2 where S is the sum of the first n terms and n is the number of terms

OpenStudy (anonymous):

so , s = 30*(72-102)/2 .

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

keep going and tell me what you get

OpenStudy (anonymous):

-450 ?

jimthompson5910 (jim_thompson5910):

i'm getting that as well

OpenStudy (anonymous):

awesome :) thanks !

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

alternatively (not the best alternative, but it's still out there) you can do this 72+66+60+54+48+42+36+30+24+18+12+6+0+(-6)+(-12)+(-18)+(-24)+(-30)+(-36)+(-42)+(-48)+(-54)+(-60)+(-66)+(-72)+(-78)+(-84)+(-90)+(-96)+(-102) = -450 Basically, generate 30 terms by repeatedly subtracting 6 from them to get 30 terms (first term is 72 and last term is -102). Then add them all up. You'll see that the answer is still -450 This is a long method, so the first method is definitely preferred.

OpenStudy (anonymous):

i tried that the first time but came up with the wrong answer ... probably a simple adding error on my part .

jimthompson5910 (jim_thompson5910):

that's probably the case I had a program help me generate all the terms which is why I was able to do it accurately...even then it gets a bit ugly